ISRAEL JOURNAL OF MATHEMATICS

TABLE OF CONTENTS, VOLUME 124, 2001

Akcoglu, M., J. Baxter, A. Bellow and R. L. Jones On restricted weak type (1,1); the discrete case

285

Amigó, J. M. Relations among sums of reciprocal powers  177

Baxter, J. See M. Akcoglu

Bellow, A. See M. Akcoglu

Berson, J., A. van den Essen and S. Maubach Derivations having divergence zero on R[X, Y]

115

Bolthausen, E. and K. Münch-Berndl Quantitative estimates of discrete harmonic measures

125

Bridson, M. R. and D. T. Wise Malnormality is undecidable in hyperbolic groups

313

Bumagina, I. Examples of amalgamated products  279

Chen, A. See K. Liu

Conca, A., E. De Negri and M. E. Rossi On the rate of points in projective spaces

253

De Negri, E. See A. Conca

Dilworth, S. J. and D. Kutzarova On the optimality of a theorem of Elton on ℓ_{1}^{n} subsystems

215

Essen, A. van den See J. Berson

Fabian, M., G. Godefroy and V. Zizler The structure of uniformly Gâteaux smooth Banach spaces

243

Fisher, D. A canonical arithmetic quotient for actions of lattices in simple groups

143

Gitik, M. and S. Shelah More on real-valued measurable cardinals and forcing with ideals

221

Gluskin, E. D., A. E. Litvak and N. Tomczak-Jaegermann An example of a complex body without symmetric projections

267

Godefroy, G. See M. Fabian

Hinrichs, A. Uniform quotient mappings of the plane with nondiscrete point inverses

203

Ivanyos, G. Deciding finiteness for matrix semigroups over function fields over finite fields: A note on a paper by Rockmore, Tan, and Beals 185 Jones, R. L. See M. Akcoglu

Kahn, J. Range of cube-indexed random walk  189

Kifer, Y., Y. Peres and B. Weiss A dimension gap for continued fractions with independent digits

61

Kutzarova, D. See S. J. Dilworth Lascar, B. et N. Lerner Propagation des singularités Gevrey pour

l’équation de Cauchy-Riemann dégénérée

299

Lerman, E. Contact cuts  77

Lerner, N. See B. Lascar Litvak, A. E. See E. D. Gluskin Liu, K. and A. Chen A class of generalized Lévy Laplacians in infinite

dimensional calculus

157

Maubach, S. See J. Berson Münch-Berndl, K. See E. Bolthausen

Neyman, A. Values of non-atomic vector measure games  1

Peres, Y. See Y. Kifer Puczyłowski, E. R. See A. Smoktunowicz

Quas, A. N. and P. B. Trow Mappings of group shifts  333

Rossi, M. E. See A. Conca Shelah, S. See M. Gitik Smoktunowicz, A. and E. R. Puczyłowski A polynomial ring that

is Jacobson radical and not nil

317

Solynin, A. Yu. and M. Vuorinen Estimates for the hyperbolic metric of the punctured plane and applications

29

Tomczak-Jaegermann, N. See E. D. Gluskin Trow, P. B. See A. N. Quas Vuorinen, M. See A. Yu. Solynin Weiss, B. See Y. Kifer Weitsman, A. Univalent harmonic mappings of annuli and a

conjecture of J. C. C. Nitsche

327

Wise, D. T. See M. R. Bridson

Xu, F. and J. Zhao Euler systems in global function fields  367
Zannier, U. Good reduction of certain covers P^{1} → P^{1}  93

Zhao, J. See F. Xu Zizler, V. See M. Fabian